OpenStudy (anonymous):

Solve xdy/dx +y = lnx.y^2

OpenStudy (shamim):

solve for what?

OpenStudy (anonymous):

\[\begin{align*}x\frac{dy}{dx}+y&=\ln x\cdot y^2\\ x\frac{1}{y^2}\frac{dy}{dx}+\frac{1}{y}&=\ln x\end{align*}\] Let \(t=y^{1-2}=\dfrac{1}{y}\), then \(\dfrac{dt}{dx}=-\dfrac{1}{y^2}\dfrac{dy}{dx}\): \[\begin{align*}-x\frac{dt}{dx}+t&=\ln x\\ \frac{dt}{dx}-\frac{1}{x}t&=-\frac{\ln x}{x}\end{align*}\] Find the integrating factor: \[\mu(x)=\exp\left(\int-\frac{dx}{x}\right)=\exp(-\ln x)=-\frac{1}{x}\] \[\begin{align*}-\frac{1}{x}\frac{dt}{dx}+\frac{1}{x^2}t&=\frac{\ln x}{x^2}\\ \frac{d}{dx}\left[-\frac{t}{x}\right]&=\frac{\ln x}{x^2}\\ -\frac{t}{x}&=\int\frac{\ln x}{x^2}~dx \end{align*}\] For the integral, integrate by parts: \[\begin{matrix}f=\ln x&&&dg=\frac{dx}{x^2}\\df=\frac{dx}{x}&&&g=-\frac{1}{x}\end{matrix}\] \[\int\frac{\ln x}{x^2}~dx=-\frac{\ln x}{x}+\int\frac{dx}{x^2}=-\frac{\ln x}{x}-\frac{1}{x}+C\] So you get \[\begin{align*}-\frac{t}{x}&=-\frac{\ln x}{x}-\frac{1}{x}+C\\\\ t&=\ln x+1-Cx \end{align*}\]

OpenStudy (anonymous):

Oh, and back-substituting again: \[\frac{1}{y}=t\]